### v1

# Low energy phonons in single crystal ZrW2O8

Abstract
ZrW2O8 is the prototypical example of a material exhibiting negative thermal expansion (NTE). As well as being among the first NTE materials to be studied in detail, it has served as a test bed for ideas for the physical mechanisms that lead to this unusual behavior. It is now widely accepted that here, and in many other framework materials exhibiting NTE, a collection of low energy phonon modes, as opposed to just one or two, are responsible for the anomalous thermal properties. However, quantitative verification and analysis of the density functional theory (DFT) calculations which underpin this proposal are still lacking. In particular, probing the low energy phonons directly throughout reciprocal space using inelastic neutron scattering, as opposed to other techniques which only probe the Brillouin zone center, is technically challenging and hence rarely done. Here we report inelastic neutron scattering measurements in a large number of Brillouin zones, achieved via the time of flight technique, over a 400 K temperature range. We find excellent agreement between DFT calculations and experimental data at low temperature, albeit with a rescaling of the calculated phonon energies, and of the lattice parameter used in the calculations compared to the measured value. However, the shifts in phonon modes due to the reduction in lattice parameter on warming, that can be predicted using DFT, are not observed. This is most likely due to counteractive anharmonic effects arising from the increased amplitude of lattice vibrations at elevated temperatures. Notwithstanding, the good agreement between experiment and DFT gives high confidence that the theory’s predictions of atomic motions for different modes is correct, and hence can be used as a reliable basis for quantitative analysis of NTE effects.
I Introduction
Negative thermal expansion (NTE) is when a material contracts rather than expands on warming. This effect can be uniform, i.e. along all crystallographic axes [1, 2, 3], or non-uniform [4, 5, 6]. It may also exist over a wide or a more limited range of temperatures [7]. ZrW2O8 is in many respects the canonical example, with its NTE having been re-discovered in the 1990s [8, 9, 10]. It is particularly notable that its NTE behavior is uniform, extant over a wide temperature range, from $\sim 10$ K to $\sim 1000$ K, and its coefficient of thermal expansion is large and negative ($\sim 10^{-5}\rm{K}^{-1}$) (see review articles ref. 11 and ref. 12 and references therein). In common with many other materials known to exhibit NTE, ZrW2O8 has a framework structure, i.e. it contains many ions that form coordinated polyhedra, most of which share corners, with large pores or spaces between the polyhedra. Indeed, it has more recently been found that intercalation of water into these pores can also result in significant anomalous lattice expansion or contraction [13]. The low temperature crystal structure of ZrW2O8 is shown in figure 1.
Figure 1: Crystal structure of ZrW2O8 with ZrO6 octahedra shaded green and WO4 tetrahedra shaded purple and blue, corresponding to tetrahedra centred on the crystallographically distinct W1 and W2 sites respectively. Note that each WO4 tetrahedron contains one unbonded oxygen ion on its vertex, pointing along a local $\langle$1 1 1$\rangle$-direction. The Zr–O–W2 bond angle is $\sim 172^{\circ}$ while the Zr–O–W1 bond angle is $\sim 154^{\circ}$. This figure was created using VESTA [14].
As a material is warmed phonon modes with gradually increasing energy are populated. If the potentials for these modes are asymmetric as a function of displacement, then thermal expansion will arise [15, 16]. Studies of the phonons in NTE materials often involve measurement of the phonon density of states (PDOS) using either Raman spectroscopy, inelastic x-ray scattering, or inelastic neutron scattering measurements of powder samples [17, 18, 19, 20, 21, 22, 23], as this is relatively straightforward to do and can be compared directly to harmonic lattice dynamics calculations such as those from density functional theory (DFT). ZrW2O8 is no exception, with the phonon density of states having been measured as a function of temperature soon after the seminal works on the NTE behavior [24]. It was found that the low energy peaks in the PDOS shift to higher energy on warming. Qualitatively, this is the opposite behavior to normal (positive thermal expansion) in which increased bond lengths on warming result in weaker interatomic potentials and hence a reduction in the energy of phonon modes.
In terms of interpretation of the PDOS measurements, and NTE more generally, early theories concerned the role of rigid unit modes (RUMs), in which Zr–O and W–O polyhedra remain undistorted but the whole structure flexes considerably along their shared vertices. As the amplitude of such modes increases, the root-mean-square distance between metal centers decreases, even though individual metal–oxygen bonds do not change in length [11]. Such modes would typically be found at the lower energy end of the phonon spectrum. However, various measurements have indicated that the picture is somewhat more complex than this. In particular, it has been found from EXAFS [25, 26, 27] that all of the polyhedra in the structure distort somewhat on warming, although the WO4 tetrahedra are noticeably more rigid than the ZrO6 octahedra. Notwithstanding, significant flex occurs along some of the Zr–O–W bonds that comprise the corner-sharing polyhedra. In this context, it would be appealing to search for a particular phonon mode that facilitates this distortion. However, numerous computational studies of the lattice dynamics in ZrW2O8 [28, 29, 30, 31] have demonstrated that there is no such single mode responsible for NTE. Rather, a large number of low energy phonon modes distributed throughout the Brillouin zone have a character that leads to NTE.
The direct measurements of phonons in ZrW2O8 (and many other NTE materials) to date have had a number of limitations. Raman spectroscopy can only measure the $\mathbf{q}=0$ Raman-active modes and thus does not provide a comprehensive picture, although the energy resolution available with this technique is usually superior to that available with neutron or x-ray spectroscopy. Neutron (or x-ray) scattering PDOS measurements are by their very nature a spherical average over reciprocal space of all possible phonon modes at a particular energy, so there is no direct sensitivity to the Q-space origin of shifts in energy with temperature. This can be ameliorated somewhat by combining and interpreting these measurements with DFT calculations. However, this approach is challenging in systems containing many atoms, and hence many phonon modes which may be more closely spaced than the energy resolution of the neutron spectrometer used for the measurements. In contrast, a measurement of a single crystal sample would allow individual modes to be isolated via measurements performed in many different Brillouin zones [32]. This would then provide a more stringent test of the DFT.
The motivation of the present study is therefore to develop a more comprehensive understanding of the low energy phonons in ZrW2O8 , by measuring in more Brillouin zones and with good energy resolution, to identify which particular individual or groups of phonon modes that are especially sensitive to temperature. Such measurements are sensitive to both the eigenvalues and eigenvectors extracted from DFT calculations, and hence pose a stringent test of the range of their validity, most notably for the low energy phonons thought to be important for NTE effects.
II Methods
ZrW2O8 crystallizes in the $P2_{1}3$ cubic space group, with a lattice parameter at room temperature of 9.16 Å[9]. The single-crystalline samples for these experiments were grown as described previously [33]. The sample used was composed of one crystal of mass 2.4 g with a mosaic of $\sim 0.3$ degrees, which is substantially less than the divergence of the incident neutron beam in the inelastic neutron scattering experiments described below, and hence unimportant for the experimental resolution. The lattice parameter was measured to be 9.176 Å  at 2 K, 9.159 Å  at 100 K, 9.1389 Å  at 300 K, and 9.125 Å  at 400 K during the alignment procedure for the inelastic scattering measurements described below. These values are consistent with those reported in the literature for powder samples [9].
II.1 Density functional theory
DFT calculations were performed using the plane-wave pseudopotential method as implemented in the CASTEP package [34]. The Perdew-Burke-Ernzehof (PBE) exchange-correlation functional [35] was used with norm-conserving pseudopotentials of the RRKJ type [36] from the Rappe and Bennett library [37], a plane-wave cutoff energy of 750 eV, and a $3\times 3\times 3$ Monkhorst-Pack [38] grid for the Brilloun-Zone of integration of electronic states. As also observed by Rimmer et al [28], the equilibrium lattice parameter at zero pressure is 9.259 Å, slightly larger than the measured low temperature lattice constant of $\sim 9.18$, typical of the PBE functional [39].
The phonon dispersion was calculated using Density Functional Perturbation Theory (DFPT) [40], with a $3\times 3\times 3$ Monkhorst-Pack grid of phonon wavevectors and Fourier interpolation of the resulting dynamical matrices, in order to sample the entire Brillouin zone. Additional calculations were performed for fixed unit cells of smaller dimension (9.23, 9.22, 9.215, 9.21 and 9.16 Å), at each of which the structure was relaxed and phonons calculated in the same way as for the fully relaxed structure. The calculated stress for each of these cell dimensions was 0.81, 1.13, 1.29, 1.45 and 3.20  GPa respectively. In the first four of these cases no soft phonon modes were observed, indicating that the computed structure remained dynamically stable. For the calculation with a lattice parameter of 9.16 Å a large number of negative energy modes were observed, indicating that for such a compressed lattice the structure becomes unstable. From this point onwards when we refer to these DFT calculations we mean those with $9.21\leq a\leq 9.23$.
When comparing the computed dispersion with that measured experimentally, an empirical scale factor is applied to the eigenvalues (energies) but not the eigenvectors (structure factors). This is determined by finding the simulation that agrees most closely with the data after such a uniform re-scale: the structure factors from the simulations with different lattice parameters are fitted to the neutron data with a single adjustable parameter (the energy scale). For ZrW2O8 the best overall agreement was found for the case of $a=9.215$and a scale factor $\sim 0.95$. This approach is used because changing the lattice parameter for the simulation does not affect the eigenvalues uniformly.
Additional calculations were performed with the same DFT parameters and a lattice parameter of 9.215 Å, but using the PBESol functional [39], which has been reported as yielding values for the bulk properties (elastic constants) of ZrW2O8  closer to the experimentally determined values than other functionals [41].
II.2 Inelastic neutron scattering
The time of flight (ToF) neutron scattering experiments were performed on the IN5 spectrometer at the Institut Laue-Langevin [42]. Data were collected with incident neutron energy $E_{i}=8.24$ meV, with the final chopper spun at 16,000 r.p.m. to give a resolution at the elastic line of 0.27 meV (full-width half-maximum, FWHM). The sample was aligned so that the $(H,H,0)$ and $(0,0,L)$ reciprocal space directions were in the horizontal plane of the instrument. The data were collected by scanning the sample orientation about the vertical axis in 0.5 degree steps over a range of 161 degrees, starting from the $(1,1,0)$-axis perpendicular to the incident neutron beam. Data were collected at temperatures of 100 K and 400 K, with these temperatures chosen in order to increase the signal via the Bose-Einstein population factor while maintaining a large temperature difference between measurements, but avoiding the structural phase transition at $\sim 425$ K [9]. Additional data were collected with the sample aligned with the $(H,0,0)$ and $(0,K,0)$ directions in the horizontal plane of the spectrometer, with the sample orientation scanned over a range of 88 degrees in 1 degree steps. These data were collected with $E_{i}=10.43$ meV and a final chopper speed of 16,000 r.p.m. giving a FWHM energy resolution at the elastic line of 0.35 meV. Data with this configuration were collected at temperatures of 2 K and 295 K.
The raw data were processed (units conversion, normalisation of intensities, etc.) using the Mantid software package [43]. Further processing of the data, and all subsequent visualisation and analysis, including convolution of calculated spectra with the instrumental resolution, were performed using the Horace software package [44]. Calculation of the inelastic neutron scattering cross-section ($S(\mathbf{Q},\omega)$) from the DFT-calculated phonons was done using the Euphonic software package’s Horace plugin [45, 46].
III Results
III.1 INS data
Figure 2 shows slices through the 4-dimensional datasets recording $S(\mathbf{Q},\omega)$, in the $\mathbf{Q}$–Energy plane, following a path through reciprocal space traversing a set of high-symmetry points. Note that although the points labelled as X, Y, and Z; and as M1, M2 and M3 are equivalent by symmetry, there are paths between them that are inequivalent by symmetry because the space group, $P2_{1}3$, is non-centrosymmetric. Data are presented centred at $\mathbf{G}=(0,0,4)$ in fig. 2 and $\mathbf{G}=(-2,-2,2)$ in fig. 3 for both $T=100$ K and $T=400$ K, and at $\mathbf{G}=(3,-3,0)$ in fig. 4 for $T=2$ K and $T=295$ K. The DFT-calculated dispersion relation with $a=9.215$, with the energy scale reduced by 5%, is overplotted (see discussion in section II.1 for further details). The calculated structure factor is also shown in panel (c) of all three figures. This set of figures illustrate many of the salient points of this work. It is clear that the dispersion and cross-section calculated using DFT provide an extremely good level of agreement with the neutron measurements, with significant overlap at most wavevectors, indicating that both the eigenvalues and eigenvectors from the DFT are accurate.
[…]
Figure 3: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2. The data here were centered on the $\mathbf{Q}=(-2,-2,2)$ Brillouin zone. Notice that the change in direction of $\mathbf{Q}$ results in the highest intensity being peaked at different high symmetry positions than for the data shown in fig. 2.
Figure 4: Slices through the $\mathbf{Q}$–Energy plane, following the same scheme as fig. 2, with the exception of the extra reciprocal lattice point on the right hand side that is labelled M3 and refers to ($\frac{1}{2},-\frac{1}{2},0)$. The data here were collected in a separate experiment and are centered on the $\mathbf{Q}=(3,-3,0)$ Brillouin zone. Data shown were collected at 2 K and 300 K, which is different to figs. 2 and 3. The different experiments also have a different (but mutually consistent) set of arbitrary units for the intensity scale. These data are notable by having intensity peaked near the G point, unlike the other two datasets shown, and extending to slightly higher energy.
At most wavevectors a change in temperature of $\sim 300$ K results in very little change in the measured dispersion and $S(\mathbf{Q},\omega)$. A broadening of the signal from most of the phonons modes, indicative of anharmonicity at elevated temperatures, is observed. One notable exception is that on warming from 2 K to 300 K a significant mode hardening of around 0.5 meV is seen on the mode between the Y and R points in fig. 4. To show this in more detail a cut at the mid-point, $\mathbf{Q}=(3.25,-2.5,-0.25)$, is shown in fig. 5. Modes along similar trajectories in the reduced Brillouin zone in the other datasets (figs. 2 and 3) do not show such a dramatic shift. However, we can see from careful inspection of the DFT-calculated dispersion at the Y point that what appears to be a single mode is actually two different ones with only a small separation in energy. The fact that the two modes appear to behave differently with temperature is a good illustration of the power of measuring in multiple Brillouin zones - with only the data centered on $(-2,-2,2)$ and $(0,0,4)$ this effect would not have been visible. A similar, although less dramatic, mode hardening occurs at the mid-point of the Z–M2 trajectory (equivalent by symmetry to Y–M1), which is visible in both fig. 2 and 3, and is shown as a cut in fig. 6.
Figure 5: Cut at the mid-point of the Y–R trajectory of fig. 4(a) and (b), i.e. at $\mathbf{Q}=(3.25,-2.5,-0.25)$, showing a significant hardening of the first phonon peak from $\sim 3.5$ meV to $\sim 4$ meV.
Figure 6: Cut at the mid-point of the Z–M2 trajectory of fig. 2(a) and (b), i.e. at $\mathbf{Q}=(0,0.25,4.5)$, showing a slight hardening of the first phonon peak from $\sim 3.4$ meV to $\sim 3.5$ meV.
Despite the overall excellent agreement between the DFT calculations and the INS measurements shown in figs. 2, 3 and 4, there are a small number of wavevectors for which there is some disagreement. These can be seen by comparing the calculated dispersion and cross-section shown in panel (c) with the data in panels (a) and (b) of all three figures. For example, we can see that for the lowest energy mode on the path M1 – Y (and M2 – Z, which is equivalent by symmetry) the calculated energy is rather higher than the measurement at 100 K, whereas the other low energy ($<4$ meV) modes along other paths agree rather closely at this temperature.
To allow a more quantitative analysis we show constant $\mathbf{Q}$ cuts along the energy axis at various high symmetry wavevectors, G, X, M and R, at different temperatures in fig. 7. In all cases the integration range along the orthogonal $q$-axes is 0.04 reciprocal lattice units. This finite integration range, necessary when plotting data from time-of-flight INS experiments, means that for modes that are highly dispersive near the wavevector of interest the width of peaks may be broader than the nominal instrumental energy resolution. This is most apparent for the single peak in the cut at R – referring to fig. 3 one can see that the dispersion is very steep until quite close to the R point.
Figure 7: Cuts along the energy axis at four high symmetry positions in reciprocal space, centered on different Brillouin zones in order to maximize the cross-section. Panel (a) shows a cut at G, panel (b) a cut at Z, panel (c) a cut at M2, and panel (d) a cut at R. Data at lower temperatures (2 K in panel (a), 100 K in panels (b)–(d)) are shown in blue whereas data at higher temperatures (300 K in panel (a), 400 K in panels (b)–(d)) are shown in red. For all cuts an estimated constant background has been subtracted, followed by correction for the Bose-Einstein population factor in order to allow easier direct comparison of data taken at different temperatures. Solid lines correspond to fits to Gaussians (the number of which is set by the number expected from the DFT calculations), to indicate the peak position.
NTE in ZrW2O8 is reflected in the phonon dispersion relation via the mode Grüneisen parameters, $\gamma_{k}$. For a phonon mode $k$ with frequency $\omega_{k}$ this is defined as
$\gamma_{k}=-\frac{V}{\omega_{k}}\frac{\partial\omega_{k}}{\partial V}$ (1)
where $V$ is the volume of the unit cell. It was by analysis of the mode Gruüneisen parameters obtained from DFT calculations [28, 29, 30, 31] that it was proposed that NTE in ZrW2O8 is dominated by a collection of low energy phonon modes of varying character. However, the way mode Grüneisen parameters are determined from DFT is rather different to what actually happens when warming and cooling a sample. In the DFT, the lattice dynamics are calculated at zero temperature in the quasi-harmonic approximation for a series of different fixed lattice parameters close to the ground state parameters determined from a minimization of the overall forces and energies. This therefore corresponds to the lattice contracting or expanding under pressure / stress, rather than warming / cooling. In reality, the warming will result in an increased amplitude of lattice vibrations, which is not accounted for in the quasi-harmonic approximation, with the consequence that anharmonic effects may become more significant at higher temperatures. This point has been discussed extensively by Mittal et al [30, 12], who have calculated this effect for a few individual modes at particular wavevectors in ZrW2O8 . Some attempt has been made to quantify the effect more broadly using Raman spectroscopy data [17, 18], which has indicated that as well as having large negative Grüneisen parameters, the lowest energy modes also have the largest anharmonicity.
There are two effects which govern the behavior of a phonon mode energy with temperature. On the one hand, in a conventional material, thermal expansion (on increasing temperature) gives rise to longer bonds between atoms which are weaker and hence give rise to lower energy phonon modes. On the other hand, as temperature increases the amplitudes of normal modes increase and intrinsic anharmonicity comes into play. This would then yield phonon modes that harden slightly with increasing temperature [16]. However, in practice in conventional materials this effect is often much smaller than the softening due to thermal expansion and so is masked. In an NTE material one might imagine that if the lattice expansion effect is opposite to normal, then both of the effects described above would add together and result in phonon mode hardening substantially with increasing temperature. Or even in a RUM scenario, in which the bonds between atoms to not change length and hence strength, the intrinsic anharmonicity would be left over to give mode hardening, albeit perhaps smaller in magnitude. Neutron measurements of the phonon density of states in ZrW2O8 [24] indeed indicate that most of the low energy modes in ZrW2O8 do indeed harden on warming. Examining fig. 7, that shows cuts along energy at several key wavevectors in different Brillouin zones, we can gain greater insight into this.
At the G-point (fig. 7(a)) at 100 K, we observe one optic mode at $5.10\pm 0.03$ meV with a FWHM of $0.41\pm 0.06$ meV. On warming to 400 K the mode is fitted to be centered at $5.12\pm 0.05$ meV, with a FWHM of $0.85\pm 0.1$ meV. Our calculations, and those of other groups, indicate that one might expect to see two modes here (see fig. 4, for example). However they are close in energy and dispersive so may not be possible to resolve and may be broadened by additional instrumental effects beyond the bare energy resolution (which is $\sim 0.23$ meV at this energy). Certainly, the peak is noticeably broadened, most likely due to thermal effects, at 400 K. The behavior of this peak contrasts with the expected softening due to the change in lattice alone of -0.2 meV (for both calculated modes), which Gupta et al [30] predict to be compensated such that the expected shifts are $-0.07$ meV and $-0.03$ meV respectively, which is significantly less than the energy resolution of our measurements  [^1] Note that the predictions of Gupta et al [30] refer to a change in temperature of 0 K to 300 K, i.e. the same difference in temperature as for our measurements. The offset is unlikely to be important, since NTE in ZrW2O8 is linear over this temperature range.. The Raman measurements of Oishi et al include this mode [18], which they find to be centered at 5.16 meV and largely independent of temperature, consistent with our finding.
At the Z-point (fig. 7(b), equivalent to the X- and Y-points) we observe three modes. At 100 K they are centered at energies of $3.50\pm 0.05$ meV, $4.43\pm 0.05$ meV and $5.28\pm 0.04$ meV. All three were found to have the same FWHM (within error) of $0.49\pm 0.05$ meV. At 400 K these peaks are centered at energies of $3.70\pm 0.10$ meV, $4.48\pm 0.09$ meV and $5.44\pm 0.13$ meV respectively. The FWHMs were found to be $0.53\pm 0.11$ meV. Gupta et al presented predictions for the first two of these modes, with a shift due to the change in lattice of -0.14 and -0.16 meV respectively, that is compensated by thermal effects to give expected shifts of 0.22 meV and 0.68 meV. The first of these predictions is consistent with our measurement, whereas the latter is not – the predicted effect of temperature (referred to as explicit anharmonicity) is larger than what is actually measured.
At the M-point (fig. 7(c), actually M2 in our notation) we observe three modes. At 100 K they are centered at energies of $3.77\pm 0.04$ meV, $4.35\pm 0.05$ meV and $5.60\pm 0.03$ meV. All three were found to have the same FWHM (within error) of $0.42\pm 0.04$ meV. At 400 K these peaks are centered at energies of $3.58\pm 0.07$ meV, $4.05\pm 0.06$ meV and $5.62\pm 0.04$ meV respectively. The FWHMs were found to be $0.34\pm 0.06$ meV. For the first two of these modes, Gupta et al predict shifts due to a change in lattice of -0.36 and -0.37 meV respectively, compensated by explicit anharmonicity to give overall predicted shifts of 0.06 and 0.18 meV. In contrast, our fits would seem to indicate that both modes soften slightly. However, inspecting fig. 3 we see that care should be taken here – it is not obvious from the colormap that these two modes soften as much as the cuts and fits seem to suggest. In fact, what appears to happen is the lowest energy mode which disperses weakly along the Z – M2 trajectory with an energy of $\sim 3.5$ meV seems to steepen slightly on warming, with the effect of bringing some of its spectral weight into the integration region of the cut at 400 K shown in fig. 7(c). Therefore, in the absence of a calculation for the shifts in energy at all $\mathbf{Q}$, we cannot quantify with much accuracy how well the measurement and prediction agree at this wavevector. The predictions suggest small shifts at about the level of our experimental resolution, which looks qualitatively reasonable.
At the R-point (fig. 7(d)) we observe one mode. At 100 K it is centered at an energy of $4.45\pm 0.02$ meV, with a FWHM of $0.56\pm 0.03$ meV. At 400 K the peak is centered at an energy of $4.43\pm 0.05$ meV with a FWHM of $0.91\pm 0.11$ meV. Here, Gupta et al predict a shift due to the change in lattice of -0.39 meV, which is complemented by a significant explicit anharmonicity of the opposite sign to those at other wavevectors, to give an overall predicted shift of -0.77 meV. The large FWHM observed at both temperatures is due to the steep dispersion of the modes near the R-point and the consequent inclusion in the cut integration region of spectral weight away from the R-point itself. Notwithstanding this, it is clear from both fig. 7(d) and fig. 3 that the mode near R does not soften as much as predicted.
III.2 DFT calculations
The DFT calculations were performed at a lattice parameter of 9.215Å, and we applied a global re-scaling of the calculated phonon energies of 95% to give best agreement with the INS measurements  [^2] This was done by doing a least squares fit of the neutron spectra with the DFT-calculated spectra, varying only the scale factor to give the best agreement. Coincidentally, the relative difference in lattice parameter between that used for the DFT and that observed experimentally at low temperature is approximately the same as the experimentally observed negative thermal expansion between 100 K and 400 K [9]. We show in fig. 8 the effect of the lattice parameter on the computed phonon spectra below 6 meV (the energy range covered in our INS experiments). As would be expected from the Grüneisen parameters reported in the literature, the low energy phonons are mostly found to harden with increasing lattice parameter. However, the scaling is not uniform, as can be seen in the inset which shows the trajectory along Z – M2 – Y. Here it is clear that the first optic mode (labelled 2 and 4 for the simulation with $a=9.259$  (the optimized lattice parameter in the DFT calculations) and $a=9.215$  respectively) is significantly more dispersive between Z and M2 for the larger lattice parameter. This also illustrates why we chose to compare the INS data with the calculations for the unit cell with $a=9.215$  rather than $a=9.259$. A global re-scaling of the latter results in a worse agreement with the measurement than the former.
Figure 8: Comparison of dispersion from DFT calculations with different lattice parameter. The reciprocal space path is the same a described in fig. 2. The solid red curve is for calculations with lattice parameter $a=9.259$(zero stress, optimized lattice parameter), and the dashed blue curve is for calculations with $a=9.215$. The inset shows a zoomed in section of the dispersion between 3 and 6 meV. The modes labelled 1, 2 and 5 correspond to calculations with $a=9.215$, and the modes labelled 3, 4 and 6 correspond to calculations with $a=9.259$. The symbols at high symmetry wavevectors correspond to the fitted peak positions for the lower temperature datasets shown in fig. 7, with diamonds for the G-point, stars for the Z-point, pentagons for the M2-point and a circle at the R-point.
One might naturally ask if further compression of the lattice in the calculations could yield yet better agreement. In fig. 9 we show calculations with $a=9.215$  and $9.210$. Here we see that certain modes, for example on the path Y – R – G, have softened considerably, such that an overall re-scale of energies would once again not yield good agreement with the INS data. What this seems to illustrate is that the calculations which agree best with the INS data are those which are on the edge of dynamical stability. Indeed, as noted in sec. II.1, when the lattice was compressed further in the DFT calculations, a significant number of soft modes appeared, consistent with this interpretation. With further compression negative energy modes appear, indicating an eventual change in the ground state structure.
Figure 9: Comparison of dispersion from DFT calculations with lattice compressed further. The reciprocal space path is the same a described in fig. 2. The solid cyan curve is for calculations with lattice parameter $a=9.210$, and the dashed blue curve is for calculations with $a=9.215$.
Figure 10: The phonon density of states (PDOS) as a function of energy computed using DFT. The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
Fig. 10 shows the partial phonon density of states (PDOS) computed from DFT, broken down into the contributions from each of the constituent atomic species. The advantage of computing this property is that it is intrinsically averaged across all of reciprocal space and hence offers some insight into the combined effect of many different phonon modes to the material’s thermal properties. The disadvantage, as already noted, is that the averaging means that information about particular phonon modes is lost.
In the energy range probed by this experiment, the PDOS has approximately equal weight from tungsten and oxygen, but very little contribution from zirconium. This indicates that the zirconium ions are much more fixed in position, somewhat characteristic of rigid-unit-modes. Further decomposition of the PDOS down to the individual atom level (see supplementary information [32]) shows that the contribution from individual tungsten ions is larger than for individual oxygen ions by about a factor two to three at low energies, but the oxygen ions are more numerous so sum up to give a comparable contribution to the PDOS.
Figure 11: The phonon density of states (PDOS) weighted by the element-specific neutron scattering factors, as a function of energy computed using DFT. The total PDOS is shown by the black curve, and the element-wise contributions are shown in red (zirconium), green (tungsten) and blue (oxygen).
Figure 12: Slices through the $\mathbf{Q}$ energy plane, following the same scheme as fig. 3. The colormaps show the neutron scattering intensity calculated using DFT, decomposed into the contributions from each individual element. Panel (a) shows the contribution from zirconium, panel (b) the contribution from tungsten, and panel (c) the contribution from oxygen. Note that the intensity scale is the same as that in fig. 3 and is the same for all three panels.
Fig. 11 shows the PDOS with neutron weighting applied (see eq. 1 in the supplementary information [32]), which allows direct comparison with the neutron structure factor measured in an experiment, such as that in ref. 24. The scattering lengths of all three constituent elements are similar in magnitude, as are their Debye-Waller factors, so the largest effect on the neutron cross-section comes from the division of all the other terms by the atomic mass. This gives significantly more weight to the oxygen modes than to tungsten and zirconium, with the consequence that the neutron scattering signal is dominated by oxygen motion. We can use a similar procedure for the single crystal plots shown in figs. 24, i.e. of element-wise decomposition of the signal. Examples are shown in fig. 12, for the $(-2,2,2)$ Brillouin zone center, i.e. equivalent to fig. 3. As with the PDOS, most of the signal arises from oxygen motions, with contributions from tungsten and zirconium around an order of magnitude smaller. However, although the neutron measurements are not directly sensitive to the motion of the heavier metal atoms, the fact that the DFT gives good agreement for the oxygen means that we can be confident that its predictions for the zirconium and tungsten motion are accurate, since each of them are surrounded by a cage of oxygen.
IV Discussion
From the results and analysis presented so far, we can see that there is generally very good agreement between the phonon dispersion relations and structure factors computed from DFT, and our inelastic neutron scattering measurements. Of course, anharmonicity is essential to describe any kind of thermal expansion, negative or positive. We have already cited extensively the predictions made by Gupta et al [30], which were based on a phenomenological theory developed for understanding librational modes at $q=0$ in the $\alpha$ phase of solid nitrogen [49]. In this approach the phonon modes are decomposed into the displacement and rotation of molecular units (analogous to the RUM approach) and the computed potential rewritten as a polynomial to $4^{\rm{th}}$ order in terms of these coordinates. This is clearly an approximation, though it should be acknowledged that it does a reasonable job, estimating correctly that in ZrW2O8 at some wavevectors the effects of implicit and explicit anharmonicity seem to approximately cancel each other.
It would seem, then, that for the low energy excitations studied here there is a cancellation of two large effects on the energy of the modes. Both arise due to anharmonicity, and given the unconventional behavior of ZrW2O8 it is clear that anharmonicity has an important role in governing the material’s thermal properties. In order to cancel the mode softening expected from the mode Grüneisen parameters, large amplitude thermal motions are required. This would naturally arise in a structure that is especially flexible, with individual atoms subject either only to weak constraints or strongly constrained only in one or two directions.
Alternative methods exist in addition to DFT of computing the phonon spectra, that include anharmonicity. These include the temperature-dependent effective potential (TDEP) method [50, 51, 52] and the self-consistent phonon (SCPH) method [53, 54]. These would in principle allow us to investigate in detail the effect of temperature and anharmonicity on the phonon modes, and hence NTE. However, the calculations to do this for a system like ZrW2O8 would be prohibitively computationally expensive, and are thus outside the scope of the present work.
Notwithstanding the aforementioned shortcomings of the DFT computations, it should be acknowledged that they perform an admirable job of reproducing many of the details of the INS measurements, both in terms of the eigenvalues (mode energies) and eigenvectors (measured neutron scattering intensities). This then gives high confidence that they can be used to give both a qualitative and quantitative commentary on the nature of the atomic displacements involved in the low energy phonon modes measured here. This was discussed extensively already by Gupta et al [30], who correlated the translational and rotational motions of quasi-rigid ZrO6 octahedra and WO4 tetrahedra, together with distortions of those polyhedra, with reported distortions found from local structure probes [55]. We can also have good confidence that the DFT results can be used as a general interpretative tool, e.g. in the flexibility analysis of NTE undertaken by Rimmer et al [28].
Molecular dynamics simulations [29] suggest that the oxygens that bridge Zr–O–W bonds are subject to large transverse vibrations. The distortions we observe from our DFT computations indicate that some modes do indeed include such motions. Because of the relative rigidity of the WO4 tetrahedra (it is well-known that the W–O bonds are shorter than the Zr–O bonds, so the WO4 tetrahedra are more tightly bonded and stiffer than the ZrO6 octahedra [56]) a consequence of such vibrations is that the lone unbonded oxygen on each tetrahedron is also subject to large vibrations [25]. Indeed, displacements of the lone oxygen would have a lower energy cost due to weaker bonding. The flexibility analysis performed by Rimmer et al [28], using DFT results similar to those reported here, showed that the low energy phonon modes responsible for most of the NTE (those with negative Grüneisen parameters) had largely fixed Zr–O and W–O bond lengths. These results together suggest that the tension effect description of NTE offers an appropriate description of the behavior of ZrW2O8 : rigid Zr–O and W–O bonds together with a large transverse vibration of oxygens that bridge Zr and W lead naturally to a reduced distance between Zr and W and hence NTE. By validating the DFT that underpins this interpretation against INS data, we are able to give greater credence to it. We also note that in the high pressure orthorhombic phase of $\gamma$-ZrW2O8 [57] the lone oxygen becomes bonded to a tungsten ion, making its motion more constrained than in the ambient pressure cubic $\alpha$-ZrW2O8 phase, which is then manifest in a significantly smaller (negative) coefficient of thermal expansion in the high pressure polymorph.
V Conclusion
In conclusion, we present a comprehensive set of INS measurements of the phonons in ZrW2O8 , and compare them to predictions from DFT. We find that we have to compress the lattice compared to the equilibrium in the DFT in order to get agreement with the neutron data, but that after doing so the agreement is excellent. This gives confidence that the overall picture of atomic motions predicted by DFT closely matches the experimental situation, so that DFT can be confidently used as a basis for additional analysis of NTE. The calculated mode Grüneisen parameters suggest that most low energy phonon modes should soften on warming as the lattice shrinks, but instead most modes either remain unchanged or harden slightly, as well as broadening somewhat. This can be understood in terms of anharmonicity arising from an increased amplitude of atomic motions associated with the phonon modes.
Acknowledgements.
We are grateful to D. Voneshen (ISIS) and R. Fair (STFC-SCD) for their advice and assistance on the use of Euphonic. We are indebted to M. Dove (QMUL) for his insightful and informative comments. Raw data are available via ref. [58]. We thank the Institut Laue-Langevin for the allocation of beamtime for the experiments reported here. Computing resources were provided by the STFC Scientific Computing Department’s SCARF cluster.